OpenStudy (anonymous):

solve.Clear the fractions first if necessary. 1/2 + 4m=3m-5/2

OpenStudy (anonymous):

multiply both sides by 2 get \[1+8m=6m-5\] then solve

OpenStudy (anonymous):

\[8m-6m=-5-1\] \[2m=-6\] \[m=-3\]